# Neural Algorithmic Reasoning with Multiple Correct Solutions

### v3

@@ Abstract
Neural Algorithmic Reasoning (NAR) aims to optimize classical algorithms. However, canonical implementations of NAR train neural networks to return only a single solution, even when there are multiple correct solutions to a problem, such as single-source shortest paths. For some applications, it is desirable to recover more than one correct solution. To that end, we give the first method for NAR with multiple solutions. We demonstrate our method on two classical algorithms: Bellman-Ford (BF) and Depth-First Search (DFS), favouring deeper insight into two algorithms over a broader survey of algorithms. This method involves generating appropriate training data as well as sampling and validating solutions from model output. Each step of our method, which can serve as a framework for neural algorithmic reasoning beyond the tasks presented in this paper, might be of independent interest to the field and our results represent the first attempt at this task in the NAR literature.
1 Introduction
@@
Classical algorithms such as Merge Sort guarantee correctness and generalize perfectly on abstract inputs [1]. Unfortunately, real problems are rarely abstract inputs. Consider pathfinding: to give traffic directions via a shortest paths algorithm, one must compress physical distance and multi-dimensional congestion data for each street into edge labels, thereby losing information. Bad edge labels mean bad directions, even despite a good algorithm. Neural Networks (NNs) avoid this problem by directly operating on high-dimensional data. Instead of compressing distance and congestion to a single path-length edge label, an NN can operate directly on distance and congestion data. The chief hurdle to NN performance is failure to generalize: common architectures are fragile to changes in input size and structure — exactly where algorithms succeed [1]. It is thus natural to wonder whether an NN might learn to generalize from an algorithm. To that end, Neural Algorithmic Reasoning trains NNs to mimic the running of classical algorithms [1]. A benefit of this approach is staying high-dimensional: no information is lost when projecting high-dimensional real data into a single abstract output for a classical algorithm [1]. Thus far, NNs have been trained to mimic algorithms which recover one solution from the space of possible correct solutions [2, 3, 4, 5, 6, 7, 8]. Ordinarily, an NN trained to perform DFS produces an array, which one can view as a distribution containing the probabilities of different solutions, from which the likeliest solution is returned as the model output. We adapt the model input to reflect multiple correct solutions and present methods of retrieving multiple solutions from the model output distribution. Using only one solution in NAR may be vulnerable to local minima in the model training. By diversifying the space of solutions returned by NAR, we may decrease the risk of solutions originating from a local minimum. For applications like cyberphysical systems, the flexibility of multiple solutions is essential. A single solution, for reasons of cost, safety, or compliance, may not be amenable to a final task.
1.1 Contributions
@@ 2 Methodology
To produce multiple solutions for a single graph, we train an NN to predict a distribution of solutions for each individual graph (solutions being individual DFS trees or BF paths, their distribution represented as distributions over child-parent pairs, as discussed in A.1.1 and A.2.1). Ordinarily, the CLRS benchmark [2] generates training data by running classical algorithms with deterministic tie-breaking. For example, the input graph in Figure 1 has two valid orders of depth-first traversal, but classical DFS will only return one based on which edge it explores first. To create our training examples, we run a classical algorithm (e.g. DFS) multiple times with randomized tiebreaking, and average the outputs (e.g. individual DFS trees) into a distribution of solutions. As we want the NN to predict a distribution, we train to minimize a distribution divergence (Kullback-Leibler: KL) [9] between the input distribution of solution and that given by the NN. Our final goal is to recover multiple solutions for an input graph; so, we develop stochastic, algorithm-specific sampling methods. Stochasticity of sampling methods means samples are likely to have multiple different solutions. For DFS, we sample the immediate parents of vertices until we reach the root vertex (which can be seen as executing DFS in reverse on a graph); our two DFS extractors, which we call Upwards and AltUpwards, differ in the manner in which they discard potential solutions (Appendix A.1.2). For BF, Beam samples a number of parents for non-source vertices according to parent-probabilities and chooses the lowest cost path arising (Appendix A.2.2). Also for BF, Greedy samples parents for each non-source vertex and chooses the parent with the lowest-cost outgoing edge for each non-source vertex (Appendix A.2.2).
@@
Figure 1: Blueprint for NAR with multiple correct solutions. For Step 2 see Figure 3. For Step 4 see Figure 4 and Appendices A.1.2 A.2.2.
3 Experiments
@@
Addressing the lack of evaluation methods for NAR with multiple solutions, we propose metrics to evaluate the validity and diversity of solutions. For a single graph, our NN predicts one solution distribution. From this distribution, we sample 5 solutions and count the number of distinct and valid solutions. For Bellman-Ford, a solution is valid if and only if it gives all shortest paths from the source vertex. For DFS, we say a solution is valid if and only if it does not contradict some necessary conditions A.2.3 (sufficient conditions not being known from tree structure alone). Figures 3 and 4 in Appendix A illustrate data generation and solution extraction on a graph. We investigate our coverage of the solution space in Appendix C. We run our experiments using the CLRS Benchmark [2], which handles data generation, model training, and model evaluation in a unified way. CLRS generates graphs of size 4, 7, 11, 13, and 16 to train our models on. In keeping with established methodology [2, 3], we test our models on graphs within and outside the range of our training sizes. For within-sample testing, we choose $n=|V(G)|=5$, as solutions on these graphs are easily manually verifiable and fast to compute at scale. For our out-of-distribution evaluation, we use the default CLRS-30 benchmark testing size $n=64$. The models are trained with default CLRS-30 benchmark model parameters (Appendix D) [2].
3.1 Results for Bellman-Ford
@@
On small graphs, ($n=5$), sampling methods extract a single correct minimum-cost path with perfect accuracy from the NN-predicted distributions (Table 1, Table 2). To validate sampling methods, we also extract solutions from the empirical distributions. Similar results for empirical and predicted distributions suggest the NN predicts a good distribution for $n=5$ graphs. The lack of variety in solutions, with sampling methods always extracting the same solution, is expected for small graphs as there is likely only one minimum-cost path. Diversity increases in the case of $n=64$ graphs, where for each method, sampling a predecessor array five times gives a new solution in nearly all cases (Table 1). For larger graphs, Table 2 shows that sampling methods successfully extract correct minimum-cost paths from the empirical distributions (denoted as $\mathcal{P}$). However, extraction is often inaccurate from the distributions of solutions predicted by our models (denoted $\tilde{\mathcal{P}}$). These results permit us to conclude that our models perform well on small graphs, but have difficulty scaling outside of their training distribution. While comparing our approach with the deterministic algorithmic frameworks of Veličković et al. [2], Ibarz et al. [3] is disadvantageous to our method due to its broader scope, it is worth noting that our BF results are markedly worse than the 99% micro-F1 score in [2, 3]. However, our results are encouraging for future work, as we have demonstrated the feasibility of training neural networks to emulate a version of Bellman-Ford that is more powerful than its deterministic variant.
@@
Table 1: Mean proportions of unique and valid solutions when sampling 5 solutions from empirical ($\mathcal{P}$) and model output ($\tilde{\mathcal{P}}$) distributions for each graph in 5 training runs. Rounded figures.
@@
Greedy ($n=5$) $0.20\pm 0.00$ $1.00\pm 0.00$ $0.20\pm 0.00$ $1.00\pm 0.00$
Beam ($n=5$) $0.20\pm 0.00$ $1.00\pm 0.00$ $0.20\pm 0.00$ $1.00\pm 0.00$
Greedy ($n=64$) $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.01$ $0.40\pm 0.19$
Beam ($n=64$) $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $0.54\pm 0.22$
@@
Upwards ($n=5$) $0.74\pm 0.04$ $0.32\pm 0.06$ $0.52\pm 0.05$ $0.16\pm 0.07$
AltUpwards ($n=5$) $0.51\pm 0.03$ $0.86\pm 0.02$ $0.73\pm 0.05$ $0.56\pm 0.14$
Upwards ($n=64$) $1.00\pm 0.00$ $0.01\pm 0.02$ $1.00\pm 0.00$ $0.00\pm 0.01$
AltUpwards ($n=64$) $1.00\pm 0.00$ $0.05\pm 0.03$ $1.00\pm 0.00$ $0.01\pm 0.02$
3.2 Results for Depth-First Search
In most cases, our sampling methods perform worse than simply using the most likely predecessor for all $v\in V(G)$, but have the advantage of recovering multiple solutions (Table 1). On smaller graphs, AltUpwards performs well on the model distributions, returning a valid DFS-tree around half of the time. We only consider necessary conditions, causing up to 40% of trees identified as "correct" to be incorrect for $n=5$ (evaluating this metric for $n=64$ is computationally prohibitive in absence of sufficient correctness conditions). DFS is known to be difficult, with NARs giving correct (single) solutions only around a quarter of the time without pointer reversal [2, 3, 4]. This contextualises our results and demonstrates the promise of our approach compared to both classical algorithms and previous work in neural algorithmic reasoning. In terms of the ability of recovering multiple solutions, Table 1 shows that for both considered graph sizes, many unique solutions are obtained, with better accuracy scores for $\mathcal{P}$ than for $\tilde{\mathcal{P}}$, particularly for large graphs, where accuracy is low in general. In fact, for the out-of-distribution testing, our NAR method for DFS performs significantly worse than the current state of the art [3]. Because our method is non-deterministic, this should not be taken as a conclusively negative result: we would expect our method for returning multiple solutions to perform worse than the results for models giving single solutions in [3], as predicting multiple solutions is a more difficult task with more noise. Furthermore, AltUpwards is generally able to recover more unique and more valid solutions than Upwards, but at a high computational cost. Table 1 shows that for large graphs, both sampling methods yield many unique solutions, but these solutions are hardly ever valid for large graphs.
Table 2: Mean accuracies (0 to 1) for all sampling methods, with standard deviations over 5 runs.

### v4

@@ Abstract.
Neural Algorithmic Reasoning (NAR) extends classical algorithms to higher dimensional data. However, canonical implementations of NAR train neural networks to return only a single solution, even when there are multiple correct solutions to a problem, such as single-source shortest paths. For some applications, it is desirable to recover more than one correct solution. To that end, we give the first method for NAR with multiple solutions. We demonstrate our method on two classical algorithms: Bellman-Ford (BF) and Depth-First Search (DFS), favouring deeper insight into two algorithms over a broader survey of algorithms. This method involves generating appropriate training data as well as sampling and validating solutions from model output. Each step of our method, which can serve as a framework for neural algorithmic reasoning beyond the tasks presented in this paper, might be of independent interest to the field and our results represent the first attempt at this task in the NAR literature.
Neural Algorithmic Reasoning, Graph Neural Networks
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Machine Learning on Graphs in the Era of Generative Artificial Intelligence; August 03–07,2025; Toronto, Canadaccs: Computing methodologies Machine learning
1. Introduction
@@
Classical algorithms like Merge Sort guarantee correctness and generalize perfectly on abstract inputs. Unfortunately, real problems are rarely abstract inputs. Consider pathfinding: to give traffic directions via a shortest paths algorithm, one must compress physical distance and congestion data for each street into single-number edge labels, thereby losing information. Bad edge labels mean bad directions, even despite a good algorithm. Neural Networks (NNs) avoid this problem by directly operating on high-dimensional data. Instead of compressing distance and congestion to a single path-length edge label, an NN can operate directly on distance and congestion data. The chief hurdle to NN performance is failure to generalize: common architectures are fragile to changes in input size and structure — exactly where algorithms succeed. It is thus natural to wonder whether an NN might learn to generalize from an algorithm. To that end, Neural Algorithmic Reasoning is the field that aims to train NNs to mimic the running of classical algorithms (Veličković and Blundell, 2021). A benefit of this approach is staying high-dimensional: no information is lost when projecting high-dimensional real data into a single abstract output for a classical algorithm (Veličković and Blundell, 2021). Thus far, NNs have been trained to mimic algorithms which recover one solution from the space of possible correct solutions (Veličković et al., 2022; Ibarz et al., 2022; Bevilacqua et al., 2023; Rodionov and Prokhorenkova, 2024; Georgiev et al., 2024; Veličković et al., 2020; Bohde et al., 2024). Ordinarily, an NN trained to perform DFS produces an array, which one can view as a distribution containing the probabilities of different solutions, from which the likeliest solution is returned as the model output. We adapt the model input to reflect multiple correct solutions and present methods of retrieving multiple solutions from the model output distribution.Using only one solution in NAR may be vulnerable to local minima in the model training. By diversifying the space of solutions returned by NAR, we may decrease the risk of solutions originating from a local minimum. For applications like cyberphysical systems, the flexibility of multiple solutions is essential. A single solution, for reasons of cost, safety, or compliance, may not be amenable to a final task.
1.1. Contributions
@@ 2. Methodology
To produce multiple solutions for a single graph, we train a NN to predict a distribution of solutions for each individual graph (solutions being individual DFS trees or BF paths, their distribution represented as distributions over child-parent pairs, as discussed in Appendices A.1.1 and A.2.1). Ordinarily, the CLRS benchmark (Veličković et al., 2022) generates training data by running classical algorithms with deterministic tie-breaking. For example, the input graph in Figure 1 has two valid orders of depth-first traversal, but classical DFS will only return one based on which edge it explores first. To create our training examples, we run a classical algorithm (e.g. DFS) multiple times with randomized tiebreaking, and average the outputs (e.g. individual DFS trees) into a distribution of solutions. As we want the NN to predict a distribution, we train to minimize a distribution divergence (Kullback-Leibler: KL) (Kullback and Leibler, 1951) between the input distribution of solution and that given by the NN. As our final goal is to recover multiple solutions for an input graph, we develop algorithm-specific stochastic sampling methods. Thus, there are two sources of randomness in our process: the translation from a deterministic algorithm to a probability distribution over predecessor-successor relationships uses randomness, as does the sampling of predecessor arrays from the model output distribution. Stochasticity of sampling methods means samples are likely to have multiple different solutions. For DFS, we sample the immediate parents of vertices until we reach the root vertex (which can be seen as executing DFS in reverse on a graph); our two DFS extractors, which we call Upwards and AltUpwards, differ in the manner in which they discard potential solutions (Appendix A.1.2). For BF, Beam samples a number of parents for non-source vertices according to parent-probabilities and chooses the lowest cost path arising (Appendix A.2.2). Also for BF, Greedy samples parents for each non-source vertex and chooses the parent with the lowest-cost outgoing edge for each non-source vertex (Appendix A.2.2).
@@
Figure 1. Blueprint for NAR with multiple correct solutions. For Step 2 see Figure 3. For Step 4 see Figure 4 and Appendices A.1.2 A.2.2. Best viewed on a screen.
Figure 2. Parent Tree encoding of a single DFS solution. The start node, 0, is its own parent, represented by the value 0 at index 0. The value 2 at index 1 indicates that vertex 2 is the parent of vertex 1. Similarly, the value 0 at index 2 indicates that vertex 0 is the parent of vertex 2.
Figure 3. Parent Probability Distribution encoding of multiple DFS solutions. The value 0.5 in row $1$ column $2$ indicates that vertex 2 is a parent of vertex 1 in 50% of correct parent trees. We call the parent probabilities obtained by repeatedly running an algorithm the true parent distribution $\mathcal{P}$. The model outputs, which we aim to be as close as possible to $\mathcal{P}$, are denoted as $\tilde{\mathcal{P}}$.
Figure 4. DFS tree extraction from a parent probability matrix: $\tilde{\mathcal{P}}$ or $\mathcal{P}$. The parent tree [0,2,2] is incorrect because there is no solution in which 2 is its own parent. [0,0,0] is incorrect, despite there being solutions in which 0 is the parent of each vertex, because there is no solution in which 0 is the parent of every vertex: if 0 is the parent of 1, 1 must be the parent of 2. Best viewed on a screen.
2.1. DFS
We consider a generalization of the deterministic DFS setting implemented in the CLRS benchmark (Veličković et al., 2022). In our setting, edge-exploration is random: from a given vertex, any of the outgoing edges may be explored first. To decrease the computational burden of verifying solutions, we require that once DFS backtracks past the root of a connected component, it restarts its search at the lowest-indexed unvisited vertex. We view this generalization is natural, modelling a setting where one conducts DFS according to a chosen sequence of starting points, without assumptions about which edges to pursue first.
Figure 5. A graph with multiple correct DFS solutions. Graph edges are drawn in black. The red tree would be the result if vertices were explored in alphabetical order, while the blue tree is one of multiple other posible valid DFS trees. Note also how the DFS-tree is encoded as a predecessor array (with the entry at position $i$ is the predecessor of the $i^{th}$ vertex in alphabetical order, which we henceforth call $\pi$.
2.1.1. Randomness in training
We generate an empirical parent distribution $\mathcal{P}$ for each graph in the test set (Figure 3) by running our randomized version of DFS 20 times (we conclude from our results in Appendix B that 20 runs provide a good approximation of the target distribution for each graph). We then train the model to minimize the KL-divergence between its predicted distribution $\tilde{\mathcal{P}}$ and $\mathcal{P}$ (Appendix A.1.1).
2.1.2. Randomness in sampling
We compare 4 methods, which we refer to as Upwards altUpwards, Argmax, and Random, for extracting DFS solutions from a parent distribution, $\tilde{\mathcal{P}}$. Our two main methods sample a parent for each vertex in a way that strives to emulate the structure of a DFS traversal while taking into account the given probability distribution. We provide detailed descriptions of our methods in Appendix A.1.2.
2.2. Bellman-Ford
2.2.1. Randomness in training
Similarly to randomising DFS, our implementation of a randomised Bellman-Ford algorithm relies on randomising the order in which paths are considered by shuffling the lists of nodes. As with DFS, we obtain a distribution of parent nodes by re-running the algorithm 20 times and computing the frequency of each parent for each node (Appendix A.2.1).
Figure 6. A graph with multiple possible minimum-cost $(source,v)$-paths. As previously, the black arrows represent the edges, with the red and blue edges showing different minimum-cost paths corresponding to their respective predecessor arrays.
2.2.2. Randomness in sampling
We use a sampling algorithm we call BF Beamsearch to sample predecessor arrays $\pi$ representing the minimum-cost $(s,v)$-paths for each vertex in the graph. This algorithm closely resembles standard Beamsearch, but it differs as it samples according to the distribution given by the model (A.2.2). To sample predecessor arrays more easily from $\mathcal{P}$ and $\tilde{\mathcal{P}}$, we also use a greedy randomised algorithm (A.2.2).
2.3. Permuting Inputs
We compare our method for generating multiple solutions by predicting solution distributions and stochastically extracting solutions, with another method: permuting the inputs to the standard CLRS benchmark model (Veličković et al., 2022). Each input graph may be encoded many times isomorphically by permuting node labels. It is thus natural to wonder whether permuting node labels will lead to a neural network extracting a new solution.
3. Experiments
@@
Addressing the lack of evaluation methods for NAR with multiple solutions, we propose metrics to evaluate the validity and diversity of solutions. We regard a solution graph $G^{\prime}$ to be valid for a randomized algorithm $A$ and a graph $G$ if $G^{\prime}$ is in the set of possible outputs of $A(G)$. We assess the diversity of solutions by considering the degree to which repeatedly sampling from the same distribution produces different solutions. For a collection of graphs and corresponding solutions, we call the fraction of correct solutions produced by a given method the accuracy of the method (referred to as graph accuracy in (Minder et al., 2023)). For a single graph, our NN predicts one solution distribution. From this distribution, we sample 5 solutions and count the number of distinct and valid solutions. For Bellman-Ford, a solution is valid if and only if it gives all shortest paths from the source vertex. For DFS, no straightforward evaluation method is known and we thus present a novel verification algorithm in Appendix A.1.3. Figures 3 and 4 illustrate data generation and solution extraction on a graph. We investigate our coverage of the solution space in Appendix C. We run our experiments using the CLRS benchmark (Veličković et al., 2022), which handles data generation, model training, and model evaluation in a unified way. The CLRS benchmark generates graphs of size 4, 7, 11, 13, and 16 to train our models on. In keeping with established methodology (Veličković et al., 2022; Ibarz et al., 2022), we test our models on graphs within and outside the range of our training sizes. For within-sample testing, we choose $n=|V(G)|=5$ and $n=16$ in order to cover the range of our training data graph sizes. For out-of-distribution evaluation, we use the default CLRS benchmark testing size $n=64$. For each graph size, we test 32 Erdős-Rényi random graphs constructed with edge probability 1/2. The models are trained with default CLRS benchmark model parameters (Appendix D) (Veličković et al., 2022).
3.1. Results for Bellman-Ford
@@
On small graphs, ($n=5$), sampling methods extract a single correct minimum-cost path with perfect accuracy from the NN-predicted distributions (Table 1, Table 2). To validate sampling methods, we also extract solutions from the empirical distributions and conclude that our sampling methods are successful across graph sizes. Similar results for empirical and predicted distributions suggest the NN predicts a good distribution for $n=5$ graphs. The lack of variety in solutions, with sampling methods always extracting the same solution, is expected for small graphs as there is likely only one minimum-cost path from the source to each target. Diversity increases for $n=16$ (Table 1), with the sampling methods producing a correct solution in all cases when sampling from the empirical distributions and in more than $90\%$ of cases when sampling from the model output distributions. In the case of $n=64$ graphs sampling a predecessor array five times gives a new solution in nearly all cases (Table 1). In Appendix C.1, we examine the diversity of the returned solution in comparison to running the randomized version of Bellman-Ford. For larger graphs, Table 2 shows that sampling methods successfully extract correct minimum-cost paths from the empirical distributions (denoted as $\mathcal{P}$). However, extraction is often inaccurate from the distributions of solutions predicted by our models (denoted $\tilde{\mathcal{P}}$). These results permit us to conclude that our models perform well on small graphs, but have difficulty scaling outside of their training distribution. Comparing our approach with the deterministic algorithmic frameworks of (Veličković et al., 2022; Ibarz et al., 2022) is disadvantageous to our method due to its broader scope. Furthermore in (Veličković et al., 2022; Ibarz et al., 2022), node accuracy is used, meaning that a given solution for a given graph and algorithm scores 50% if half of its parent-successor relationships are the same as in the (single) correct solution for the graph. There is no such optimal solution that we could compare a given solution against in our case and we thus use graph accuracy (see (Minder et al., 2023) and Appendix E for further discussion of the metrics). It is important to note that graph accuracy is much stricter than node accuracy: a method with node accuracy of 92% would be expected to score only 1% graph accuracy on a $n=64$ graph. Therefore, it is difficult to compare our results to previous work in NAR. However, our results are encouraging for future work, as we have demonstrated the feasibility of training neural networks to emulate a version of Bellman-Ford that is more powerful than its deterministic variant.
@@
Table 1. Mean proportions of unique and valid solutions when sampling 5 solutions from empirical ($\mathcal{P}$) and model output ($\tilde{\mathcal{P}}$) distributions for each graph in 5 training runs. Rounded figures.
@@
Greedy (n=5) $0.20\pm 0.0$ $1.00\pm 0.0$ $0.20\pm 0.0$ $1.00\pm 0.0$
Beam (n=5) $0.20\pm 0.0$ $1.00\pm 0.0$ $0.20\pm 0.0$ $1.00\pm 0.0$
Greedy (n=16) $0.35\pm 0.0$ $1.00\pm 0.0$ $0.40\pm 0.1$ $0.92\pm 0.0$
Beam (n=16) $0.34\pm 0.0$ $1.00\pm 0.0$ $0.39\pm 0.0$ $0.90\pm 0.0$
Greedy (n=64) $1.00\pm 0.0$ $1.00\pm 0.0$ $1.00\pm 0.0$ $0.40\pm 0.2$
Beam (n=64) $1.00\pm 0.0$ $1.00\pm 0.0$ $1.00\pm 0.0$ $0.54\pm 0.2$
@@
Upwards (n=5) $0.73\pm 0.0$ $0.36\pm 0.0$ $0.45\pm 0.0$ $0.19\pm 0.0$
AltUpwards (n=5) $0.60\pm 0.0$ $0.80\pm 0.0$ $0.66\pm 0.0$ $0.77\pm 0.0$
Upwards (n=16) $1.00\pm 0.0$ $0.10\pm 0.0$ $1.00\pm 0.0$ $0.01\pm 0.0$
AltUpwards (n=16) $1.00\pm 0.0$ $0.18\pm 0.0$ $1.00\pm 0.0$ $0.14\pm 0.0$
Upwards (n=64) $1.00\pm 0.0$ $0.02\pm 0.0$ $1.00\pm 0.0$ $0.01\pm 0.0$
AltUpwards (n=64) $1.00\pm 0.0$ $0.05\pm 0.0$ $1.00\pm 0.0$ $0.03\pm 0.0$
3.2. Results for Depth-First Search
Our results indicate that even on small graphs, our methods do not achieve a good accuracy (Table 2). Without modifications, such as hint pointer reversal (Bevilacqua et al., 2023), DFS is a difficult problem for neural algorithmic reasoners (Ibarz et al., 2022; Veličković et al., 2022), and our results point towards the additional difficulty of learning to execute DFS when considering multiple solutions. As in the discussion of our results for Bellman-Ford, we point out that no suitable comparison exists with previous work. In (Ibarz et al., 2022), it is indicated the node-level micro-F1 score for DFS does not exceed 60% for $n=64$ without hint pointer reversal (Bevilacqua et al., 2023) (these values being an improvement over the previous state of the art), making graph-level accuracies significantly above 0% unlikely. Thus, while our reported accuracies are low, they are likely to be in line with previous work in the NAR, with the additional difficulty of retrieving multiple solutions. Considering the accuracy results and the fact that even on the solution distributions, our sampling methods do not produce correct solutions, we hypothesize that the encoding of multiple solutions into a probability distribution, as presented above, leads to a lack of differentiation between different correct solutions and thus to difficulties retrieving correct solutions. Further evidence for this hypothesis is the observation that there are fewer correct solutions to the single-source shortest path problem on one of the random weighted graphs we sample than there are valid depth-first traversals of that same graph. This is the case because for a given unweighted graph $G$, every path from the root could be included in a depth-first traversal, but only the shortest paths from the root to the vertices can be part of the output of the randomized Bellman-Ford algorithm on a given weighted version of $G$.
Table 2. Mean accuracies (0 to 1) for all sampling methods, with standard deviations over 5 runs.